I am a parent trying to help my son figure out how to solve the following:
Bill is twice as old as his brother Dan. In 7 years Bill will be only one and one half as old as Dan. How old is Bill now?
I would like the details on solving so I can teach him properly.
Thank you.

Answer by longjonsilver(2297)

First define the things you are asked for:

Let Bill be x years NOW
Let Dan be y years NOW.

so, we know x = 2y --eqn1

Also, in 7 years time, Bill will be x+7 and Dan will be y+7. At this point, (x+7) = (3/2)(y+7) --eqn2

eqn2 simplifies to 2x+14 = 3y+21
--> 2x - 3y = 7

Now, from eqn1, x = 2y, so sub this into the above version of eqn2, which becomes: 2(2y) - 3y = 7, which becomes

4y - 3y = 7
y = 7

Hence x=2(7) --> x=14

So, Bill is 14 years old now.

check:
in 7 years, we have ages 14 and 21, where 21 is "one and one-half" as much as 14.

Hope this is clear.

jon.